# Transition Constrained Bayesian Optimization via Markov Decision Processes

### v2

@@ C.2 Upper-bounding the objective: Eliminating $\mathbb{E}_{f}$
$\min_{\mathbf{X}_{\text{next}}}\mathbb{E}_{f}\left[\sup_{z\in\mathcal{Z}\setminus\{x_{f}^{\star}\}}\frac{k_{\mathbf{X}\cup\mathbf{X}_{\text{new}}}(z,x_{f}^{\star})}{(f(z)-f(x_{f}^{\star}))^{2}}\right]$$\leq$ $\mathbb{E}_{f}\left[\frac{1}{\operatorname{gap}(f)}\right]\min_{\mathbf{X}_{\text{next}}}\sup_{z\in\mathcal{Z}\setminus\{x_{f}^{\star}\}}k_{\mathbf{X}\cup\mathbf{X}_{\text{new}}}(z,x_{f}^{\star})$
$\leq$ $\mathbb{E}_{f}\left[\frac{1}{\operatorname{gap}(f)}\right]\min_{\mathbf{X}_{\text{next}}}\sup_{z,z^{\prime}\in\mathcal{Z}}\operatorname{Var}\left[f(z)-f(z^{\prime})|\mathbf{X}_{t}\cup\mathbf{X}_{\text{new}}\right].$
As the constant in front of the objective does not influence the optimization problem, we do not need to consider it when defining the utility. Surprisingly, this objective coincides with the objective from Fiez et al. [12] which has been derived as lower bound to the best-arm identification problem (maximum identification) with linear bandits. Their perspective is however slightly different as they try to minimize $T$ for a fixed $\delta$ failure rate. Perhaps it should not be surprising that the dual variant, consider here, for fixed $T$ and trying to minimize the failure rate leads to the same decision for large $T$ when $\log(b_{z})$ can be neglected.
@@ C.4 Theory: convergence to the optimal policy
where in the above $\hat{d}_{t}\oplus d$ represents the weighted sum as in the main Algorithm 1 that scales them properly according to $t$ and $T$. Notice that in contrast to our objective, it does not take the maximum over $z^{\prime}\in\mathcal{Z}$, but fixes it to the value $x_{f}^{\star}$ that the hypothetical algorithm has privileged access to. To eliminate the cumbersome notation, we will refer to the objectives as $\mathcal{U}(d|\mathcal{Z}_{t},\mathcal{Z}_{t})$ as the objective used by our algorithm (real execution) and $\mathcal{U}(d|\mathcal{Z}_{t},\{x_{f}^{\star}\})$, as the objective that the privileged algorithm is optimizing which serves as theoretical baseline.
The visitation of $d^{\star}_{t}$ represents the best possible investment of the resources (of the size $T-t$) to execute at time $t$ had we known the $x^{\star}_{f}$ instead of only $\mathcal{Z}_{t}$. This is interpreted as if the modeller knows $x^{\star}_{f}$, and sets up an optimal curriculum that is being shown to an observer in order to convince him/her of that $x^{\star}_{f}$ is the optimal value. He or she is using statistical testing to elucidate it from execution of the policy. Like the algorithm, the optimal policy changes along the optimization procedure due to changes in $\mathcal{Z}_{t}$. Hence, our goal is to show that we are closely tracking the performance of these optimal policies in time $t$, and eventually there is little difference between our sequence of executed policies (visitations) $\hat{d}_{t}$ and the algorithm optimal $d^{\star}_{t}$.
@@
The proof is based on the proof of Frank-Wolfe convergence that appears Appendix B.4 in Thm. 3. in Mutny et al. [39].
@@
First notice that $\frac{1}{T-1}\sum_{t=1}^{T-1}\epsilon_{t}\leq\frac{B}{\sqrt{T}}\log(1/\delta)$ by Lemma in Mutny et al. [39] due to $\epsilon_{t}$ being martingale difference sequence. The other term is the sum on $\frac{1}{T-1}\sum_{t=1}C_{t}$ which appears in the main result. The sum on $\sum_{t=1}^{T-1}\frac{1}{1+t}\leq L\frac{\log T}{T}$. The rest is eliminated by the reccurence of the terms, and using that $\mathcal{U}(d|\mathcal{Z}_{t},\{x_{f}^{\star}\}\leq\mathcal{U}(d|\mathcal{Z}_{t-1},\{x_{f}^{\star}\})$ for any $d$. This is due to set $\mathcal{Z}_{t}$ decreasing over time. We report the result in asymptotic notation as function of $T$ and $\log(1/\delta)$. ∎
@@ Appendix D Objective reformulation and linearization
For the main objective we try to optimize over a subset of $T$ trajectories $\mathbf{X}\subset\mathcal{X}^{H}$. Let be $\mathcal{X}^{H}$ be the set of sequences of inputs $\tau=(x_{1},...,x_{H})$ where they consist of states in the search space $\mathcal{X}$. Furthermore, assume there exists, in the deterministic environment, a constraint such that $x_{h+1}\in\mathcal{C}(x_{h})$ for all $h=1,...,H-1$. Then we seek to find the set $\mathbf{X}_{*}$, consisting of $T$ trajectories (possibly repeated), such that we solve the constrained optimization problem:
$\mathbf{X}_{*}=\operatorname*{arg\,min}_{\mathbf{X}\in\mathcal{X}^{H}}\max_{z,z^{\prime}\in\mathcal{Z}}\text{Var}[f(z)-f(z^{\prime})|\mathbf{X}]\quad\text{ s.t. }\quad x_{h+1}\in\mathcal{C}(x_{h})\quad\forall t=1,...,h-1$ (21)
@@
Our goal is to show that optimization over sequences can be simplified to state-action visitations as in Mutny et al. [39]. For this, we require that the objective depends additively involving terms $x,a$ separately. We formalize this in the next result. In order to prove the result, we utilize the theory of reproducing kernel Hilbert spaces [10].
@@
Let $\mathbf{X}$ be a collection of $t$ trajectories of length $H$. Assuming that $f\sim\mathcal{GP}(0,k)$. Assuming that $k$ has Mercer decomposition as $k(x,y)=\lambda_{k}\phi_{k}(x)\phi_{k}(y)$.
@@
Let $d_{\mathbf{X}}$ be the visitation of the states-action in the trajectories in $\mathbf{X}$, as $d_{\mathbf{X}}=\frac{1}{HT}\sum_{t=1}^{T}\sum_{x,a\in\tau_{t}}\delta_{x,a}$, where the $\delta$ represent delta function supported on $x,a$. Then optimization of the objective Eq. (21) can be rewritten as:
$U(d_{\mathbf{X}})=\max_{z,z^{\prime}\in\mathcal{Z}}||\Phi(z)-\Phi(z^{\prime})||_{\mathbf{V}(d_{\mathbf{X}})^{-1}}^{2},$
where $\mathbf{V}(d)=\sum_{i}\sum_{x,a\in\tau_{i}}d(x,a)\Phi(x)\Phi(x)^{\top}+\mathbf{I}/(TH)$ is a operator $\mathbf{V}(d):\mathcal{H}_{k}\rightarrow\mathcal{H}_{k}$, the norm is RKHS norm, and $\Phi(z)_{k}=\phi_{k}(z)$.
@@
Let $k_{t}(z)=\Phi(\mathbf{X})\phi(z)$. Utilizing $k(z,z^{\prime})=\Phi(z)^{\top}\Phi(z)$ (RKHS inner product) with the Mercer decomposition and the matrix inversion lemma, the above can be written as using $\mathbf{V}=\sum_{t=1}^{T}\sum{x\in\tau_{t}}\Phi(x)\Phi(x)^{\top}+\sigma^{2}\mathbf{I}_{\mathcal{H}}.$
@@
$=$ $(\Phi(z)-\Phi(z^{\prime}))\left(\frac{TH}{TH}\sum_{t=1}^{T}\sum_{x\in\mathcal{X}}\#(x\in\tau_{t})\Phi(x)\Phi(x)^{\top}+\mathbf{I}_{\mathcal{H}_{k}}\sigma^{2}\right)^{-1}(\Phi(z)-\Phi(z^{\prime}))$
$=$ $TH(\Phi(z)-\Phi(z^{\prime}))\left(TH\sum_{x\in\mathcal{X}}d(\mathbf{X})\Phi(x)\Phi(x)^{\top}+\frac{\mathbf{I}_{\mathcal{H}_{k}}}{TH}\sigma^{2}\right)^{-1}(\Phi(z)-\Phi(z^{\prime}))$
@@
In other words, optimization over trajectories and optimization over $d_{\pi}\in\mathcal{D}$ is not equivalent. The latter is merely a continuous relaxation of discrete optimization problems to the space of Markov policies. It is in line with the classical relaxation approach addressed in experiment design literature with a rich history, e.g., Chaloner and Verdinelli [7]. For introductory texts on the topic, consider Pukelsheim [42] for the statistical perspective and Boyd and Vandenberghe [5] for the optimization perspective. However, as Mutny et al. [39] points out, optimization of this objective does optimize objective as Eq. (21). In other words, by optimizing the relaxation with a larger budget of trajectories or horizons, we are able to decrease Eq. (21) as well.
For completeness, we state the auxiliary lemma.
@@ D.1 Objective formulation for general kernel methods
We now show how to write the objective in terms of the general kernel matrix instead of relying on finite dimensional embeddings. To do so we use the following well known Woodbury lemma [60]:
Lemma D.3 (Woodbury Matrix Identity).
We now work backwards from (23), and first write the objective in terms of features of arbitrarily large size. We define $\tilde{\sigma}^{2}=\sigma^{2}/TH$, a diagonal matrix that describes the state-action distribution $\mathbf{D}=\text{diag}(\{d_{x}:x\in\mathcal{X}\})$ of the size $|\mathcal{X}|\times|\mathcal{X}|$, and $\Phi(\mathcal{X})$ which corresponds to the unique (possibly infinite-dimensional) embeddings of each element in $\mathcal{X}$, ordered in the same way as $\mathbf{D}$.
@@
which allows us to calculate the objective for general kernel methods at the cost of an $|\mathcal{X}|\times|\mathcal{X}|$ inversion. Upon identifying the $z,z^{\prime}$ that maximize the above, we can use them in an optimization procedure. This holds irrespective of whether the state space is discrete or continuous. In continuous settings however, we again require a parametrization of the infinite dimensional probability distribution by some finite means.
@@ D.2 Linearizing the objective
$\Sigma_{\theta,d}$ $=\left(\sum_{x\in\mathcal{X}}\Phi(x)\Phi(x)^{T}d(x)+\lambda\sigma^{2}I\right)^{-1}$
@@ E.2 Benchmark Details
$k_{rbf}(x,x^{\prime})=\sigma_{rbf}^{2}\exp\left(-\frac{(x-x^{\prime})^{2}}{2\ell_{rbf}}\right)$

### v3

@@ C.1 Derivation of the Bayesian utility
$-\mathbb{E}_{f}\left[\sup_{z\in\mathcal{Z}\setminus\{x_{f}^{\star}\}}\frac{(f(z)-f(x_{f}^{\star}))^{2}}{k_{\mathbf{X}\cup\mathbf{X}_{\text{new}}}(z,x_{f}^{\star})}\right]$$\leq$ $-\sup_{z\in\mathcal{Z}\setminus\{x_{f}^{\star}\}}\frac{\mathbb{E}_{f}\left[\operatorname{gap}(f)\right]}{k_{\mathbf{X}\cup\mathbf{X}_{\text{new}}}(z,x_{f}^{\star})}$
$\leq$ $-\mathbb{E}_{f}\left[\operatorname{gap}(f)\right]\sup_{z\in\mathcal{Z}\setminus\{x_{f}^{\star}\}}\frac{1}{\operatorname{Var}\left[f(z)-f(x_{f}^{\star})|\mathbf{X}_{t}\cup\mathbf{X}_{\text{new}}\right]}.$
As the constant in front of the objective does not influence the optimization problem, we do not need to consider it when defining the utility. Furthermore, in order to minimise the probability of an error we can just minimise the variance in the denominator instead (since $\arg\min_{x}-g(x)$ is equivalent to $\arg\min_{x}\frac{1}{g(x)}$ when $g(x)>0$). However, the non-trivial distribution of $f(x^{\star})$ [29] renders the utility intractable; therefore we employ a simple and tractable upper bound on the objective by minimising the uncertainty among all pairs in $\mathcal{Z}$:
Surprisingly, this objective coincides with the objective from Fiez et al. [30] which has been derived as lower bound to the best-arm identification problem (maximum identification) with linear bandits. Their perspective is however slightly different as they try to minimize $T$ for a fixed $\delta$ failure rate. Perhaps it should not be surprising that the dual variant, consider here, for fixed $T$ and trying to minimize the failure rate leads to the same decision for large $T$ when $\log(b_{z})$ can be neglected.
@@ C.4 Theory: convergence to the optimal policy
where in the above $\hat{d}_{t}\oplus d$ represents the weighted sum as in the main Algorithm 1 that scales the distributions properly according to $t$ and $T$, so to make the sum of them a valid distribution. Notice that in contrast to our objective, it does not take the maximum over $z^{\prime}\in\mathcal{Z}$, but fixes it to the value $x_{f}^{\star}$ that the hypothetical algorithm has privileged access to. To eliminate the cumbersome notation, we will refer to the objectives as $\mathcal{U}(d|\mathcal{Z}_{t},\mathcal{Z}_{t})$ as the objective used by our algorithm (real execution) and $\mathcal{U}(d|\mathcal{Z}_{t},\{x_{f}^{\star}\})$, as the objective that the privileged algorithm is optimizing which serves as theoretical baseline.
The visitation of $d^{\star}_{t}$ represents the best possible investment of the resources (of the size $T-t$) to execute at time $t$ had we known the $x^{\star}_{f}$ instead of only $\mathcal{Z}_{t}$. This is interpreted as if the modeler knows $x^{\star}_{f}$, and sets up an optimal curriculum that is being shown to an observer in order to convince him/her of that $x^{\star}_{f}$ is the optimal value. He or she is using statistical testing to elucidate it from execution of the policy. Like the algorithm, the optimal policy changes along the optimization procedure due to changes in $\mathcal{Z}_{t}$. Hence, our goal is to show that we are closely tracking the performance of these optimal policies in time $t$, and eventually there is little difference between our sequence of executed policies (visitations) $\hat{d}_{t}$ and the algorithm optimal $d^{\star}_{t}$.
@@
The proof is based on the proof of Frank-Wolfe convergence that appears Appendix B.4 in Thm. 3. in Mutný et al. [17].
@@
First notice that $\frac{1}{T-1}\sum_{t=1}^{T-1}\epsilon_{t}\leq\frac{B}{\sqrt{T}}\log(1/\delta)$ by Lemma in Mutný et al. [17] due to $\epsilon_{t}$ being martingale difference sequence. The other term is the sum on $\frac{1}{T-1}\sum_{t=1}C_{t}$ which appears in the main result. The sum on $\sum_{t=1}^{T-1}\frac{1}{1+t}\leq L\frac{\log T}{T}$. The rest is eliminated by the reccurence of the terms, and using that $\mathcal{U}(d|\mathcal{Z}_{t},\{x_{f}^{\star}\}\leq\mathcal{U}(d|\mathcal{Z}_{t-1},\{x_{f}^{\star}\})$ for any $d$. This is due to set $\mathcal{Z}_{t}$ decreasing over time. We report the result in asymptotic notation as function of $T$ and $\log(1/\delta)$. ∎
@@ Appendix D Objective reformulation and linearization
For the main objective we try to optimize over a subset of $T$ trajectories $\mathbf{X}=\{\tau_{i}\in\mathcal{X}^{H}\}_{i=1}^{T}$. Let $\mathcal{X}^{H}$ be the set of sequences of inputs $\tau=(x_{1},...,x_{H})$ where they consist of states in the search space $\mathcal{X}$. Furthermore, assume there exists, in the deterministic environment, a constraint such that $x_{h+1}\in\mathcal{C}(x_{h})$ for all $h=1,...,H-1$. Then we seek to find the set $\mathbf{X}_{*}$, consisting of $T$ trajectories (possibly repeated), such that we solve the constrained optimization problem:
$\mathbf{X}_{*}=\operatorname*{arg\,min}_{\mathbf{X}\in\mathcal{X}^{TH}}\max_{z,z^{\prime}\in\mathcal{Z}}\text{Var}[f(z)-f(z^{\prime})|\mathbf{X}]\quad\text{s.t. }\quad x_{h+1}\in\mathcal{C}(x_{h})\quad\forall t=1,...,h-1$ (23)
@@
Our goal is to show that optimization over sequences can be simplified to state-action visitations as in Mutný et al. [17]. For this, we require that the objective depends additively involving terms $x,a$ separately. We formalize this in the next result. In order to prove the result, we utilize the theory of reproducing kernel Hilbert spaces [72].
@@
Let $\mathbf{X}$ be a collection of $t$ trajectories of length $H$. Assuming that $f\sim\mathcal{GP}(0,k)$. Assuming that $k$ has Mercer decomposition as $k(x,y)=\sum_{k}\lambda_{k}\phi_{k}(x)\phi_{k}(y)$.
@@
Let $d_{\mathbf{X}}$ be the visitation of the states-action in the trajectories in $\mathbf{X}$, as $d_{\mathbf{X}}=\frac{1}{TH}\sum_{t=1}^{T}\sum_{x,a\in\tau_{t}}\delta_{x,a}$, where the $\delta_{x,a}$ represent delta function supported on $x,a$. Then optimization of the objective Eq. (23) can be rewritten as:
$U(d_{\mathbf{X}})=\frac{1}{TH}\max_{z,z^{\prime}\in\mathcal{Z}}||\Phi(z)-\Phi(z^{\prime})||_{\mathbf{V}(d_{\mathbf{X}})^{-1}}^{2},$
where $\mathbf{V}(d)=\sum_{i}\sum_{x,a\in\tau_{i}}d(x,a)\Phi(x)\Phi(x)^{\top}+\mathbf{I}\sigma^{2}/(TH)$ is a operator $\mathbf{V}(d):\mathcal{H}_{k}\rightarrow\mathcal{H}_{k}$, the norm is RKHS norm, and $\Phi(z)_{k}=\phi_{k}(z)$.
@@
Utilizing $k(z,z^{\prime})=\Phi(z)^{\top}\Phi(z)$ (RKHS inner product) with the Mercer decomposition we know that $k_{t}(z)=\Phi(\mathbf{X})\phi(z)$. Applying the matrix inversion lemma, the above can be written as using $\mathbf{V}=\sum_{t=1}^{T}\sum_{x\in\tau_{t}}\Phi(x)\Phi(x)^{\top}+\sigma^{2}\mathbf{I}_{\mathcal{H}}.$
@@
$=$ $(\Phi(z)-\Phi(z^{\prime}))\left(\frac{TH}{TH}\sum_{t=1}^{T}\sum_{x\in\mathcal{X}}\#(x\in\tau_{t})\Phi(x)\Phi(x)^{\top}+\sigma^{2}\mathbf{I}_{\mathcal{H}_{k}}\right)^{-1}(\Phi(z)-\Phi(z^{\prime}))$
to arrive at:
$\text{Var}\left[f(z)-f(z^{\prime})|\mathbf{X}\right]=\frac{(\Phi(z)-\Phi(z^{\prime}))}{TH}\left(\sum_{x\in\mathcal{X}}d(\mathbf{X})\Phi(x)\Phi(x)^{\top}+\frac{\sigma^{2}}{TH}\mathbf{I}_{\mathcal{H}_{k}}\right)^{-1}(\Phi(z)-\Phi(z^{\prime}))$ (26)
@@
In other words, optimization over trajectories and optimization over $d_{\pi}\in\mathcal{D}$ is not equivalent. The latter is merely a continuous relaxation of discrete optimization problems to the space of Markov policies. It is in line with the classical relaxation approach addressed in experiment design literature with a rich history, e.g., Chaloner and Verdinelli [73]. For introductory texts on the topic, consider Pukelsheim [74] for the statistical perspective and Boyd and Vandenberghe [75] for the optimization perspective. However, as Mutný et al. [17] points out, reducing the relaxed objective does reduce the objective as Eq. (23) as well. In other words, by optimizing the relaxation with a larger budget of trajectories or horizons, we are able to decrease Eq. (23) as well.
For completeness, we state the auxiliary lemma. We make use of the Sherman-Morrison-Woodbury (SMW) formula, [76]:
Lemma D.2 (Sherman-Morrison-Woodbury (SMW)).
to show the following.
@@ D.1 Objective formulation for general kernel methods
The previous discussion also allows us to write the objective in terms of the general kernel matrix instead of relying on finite dimensional embeddings. The modification is very similar and relies again on Sherman-Mirrison-Woodbury lemma.
We now work backwards from (26), and first write the objective in terms of features of arbitrarily large size. Using the shorthand, $\tilde{\sigma}^{2}=\sigma^{2}/TH$, let us define a diagonal matrix that describes the state-action distribution $\mathbf{D}=\text{diag}(\{d_{x}:x\in\mathcal{X}\})$ of the size $|\mathcal{X}|\times|\mathcal{X}|$, and $\Phi(\mathcal{X})$ which corresponds to the unique (possibly infinite-dimensional) embeddings of each element in $\mathcal{X}$, ordered in the same way as $\mathbf{D}$.
@@
which allows us to calculate the objective for general kernel methods at the cost of an $|\mathcal{X}|\times|\mathcal{X}|$ inversion. Upon identifying the $z,z^{\prime}$ that maximize the above, we can use them in an optimization procedure. This holds irrespective of whether the state space is discrete or continuous. In continuous settings however, we again require a parametrization of the infinite dimensional probability distribution by some finite means such as claiming that $\mathbf{D}_{\theta}$ contains some finite dimensional simplicity. This is what we do with the linear system example in Section 4.2.
@@ D.2 Linearizing the objective
$\Sigma_{\theta,d}$ $=\left(\sum_{x\in\mathcal{X}}\Phi(x)\Phi(x)^{T}d(x)+\sigma^{2}I\right)^{-1}$
@@
In the definition above we dropped the constant pre-factors since they do not influence the maximizer of the gradient as they are related by a constant multiplicative factor.
@@ E.2 Benchmark Details
$k_{rbf}(x,x^{\prime})=\sigma_{rbf}^{2}\exp\left(-\frac{||x-x^{\prime}||_{2}^{2}}{2\ell_{rbf}}\right)$